OpenStudy (anonymous):

y^2 + 36x +8y - 92 = 0 help! what are the vertices, intercept, foci, eccentricity and directrices. Help please

OpenStudy (campbell_st):

so its in the form (y - k)^2 = 4a(x-h) so y^2 +8y = 36x +92 complete the square for y (y + 4)^2 = 36x + 92 + 16 (y + 4)^2 = 36(x + 3) so the vertex is at (-3, -4) a = 9 the directrix is x = -12 the parabola is concave right focus is (6, -4)

OpenStudy (anonymous):

how do i get the intercepts?

OpenStudy (moongazer):

There is NO eccentricity and intercepts for parabola. Eccentricity and intercepts are just for Ellipse and Hyperbola.

OpenStudy (moongazer):

y^2 + 36x +8y - 92 = 0 is the gen. equation of parabola

OpenStudy (anonymous):

why is that x = -12?? could you help me.

OpenStudy (moongazer):